OpenStudy (anonymous):

3x-2(4+1/3x)=-1. Solve for x.

Directrix (directrix):

3

OpenStudy (anonymous):

3x-2(4+1/3x)=-1 3x-8-2/3x= -1 3x - 2/3x= 7 9/3x - 2/3x=7 7/3x= 7 x= (7)*3/7 = 3

Directrix (directrix):

3x-2(4+1/3x)=-1